Johansen Company had $210,000 of operating income in 2020 when the selling price per unit was $169, the variable costs per unit were $119, and the fixed costs were $415,000. Management expects per-unit data and total fixed costs to remain the same in 2021. The president of Johansen Company is under pressure from shareholders to increase operating income by $65,000 in 2021.

a. Calculate the number of units sold in 2020.

b. Calculate the number of units that would have to be sold in 2021 to reach the shareholders’ desired profit level. (Round answer to 0 decimal places, e.g. 125.)
Answer #1

A) Let units sold be x

(Contribution margin per unit * Units sold) - Fixed cost = operating income

[(169-119)*x] - 415,000 = 210,000

50x - 415,000 = 210,000

x = 12,500

Units sold = 12,500

Desired sales units = (Fixed costs +Target profit) /Contribution margin per unit

= (415,000+210,000+65,000)/(169-116)

= 690,000/50

= 13,800 units